OpenStudy (melissa_something):

Really hard & annoying math: Find and simplify the different quotient of f:

OpenStudy (melissa_something):

f(x+h)-f(x) --------- where \[h \neq 0 \] h

OpenStudy (melissa_something):

Oops, the h is supposed to be under the fraction

OpenStudy (melissa_something):

f(x)= 1/8

OpenStudy (zarkon):

if f(x)=1/8 then f(x+h)=1/8 and f(x+h)-f(x)=0

OpenStudy (melissa_something):

Is that like something obvious? I never understood this. @Zarkon

OpenStudy (anonymous):

if h is tending to zero then this is the derivative of F(x)

OpenStudy (melissa_something):

Hey, what do you mean tending to 0? I haven't learned the derivatives! @cooldude17

OpenStudy (anonymous):